Fipronil is a broad-use insecticide commonly used in flea control products for pets. The experimentally determined logP (or logKow) is 4. If this value was determined experimentally, what would you expect the concentration ratio to be for fipronil in octanol: fipronil in water?

A. 1,000:1

B. 1:10,000

C. 100:1

D. 10,000:1

E. 4:1
Answer #1

log P (log Kow) = 4/1

log P /log Kow = Unionised drug in organic/octanol layer /unionised drug in aqueous layer = 4

Then experimentally concentration ratio to be for fipronil in octanol: fipronil in water = 10P/logKOW = 104 = 10000/1 = 10000 :1 (Option D) is correct answer